Solve for x: square root 5x-6 = x-4
Answer by edjones(8007)

sqrt(5x-6)=x-4
5x-6=(x-4)^2 square each side
5x-6=x^2-8x+16
x^2-13x+22=0
(x-11)(x-2)=0
x=11, x=2
.
Ed